# Enhancing Photometric Redshift Catalogs Through Color-Space Analysis: Application to KiDS-Bright Galaxies

### v1

@@ 4.2 Selection of a clean sample using SOM
In the previous section, we quantified the relative completeness of the GAMA spectroscopic data with respect to the photometric KiDS-Bright as a function of several galaxy properties. We observed that the GAMA completeness quickly drops at the faintest and highest-$z$ end, while no significant deterioration is present with regard to colors or stellar mass. As GAMA was used to train the KiDS-Bright photo-$z$, here we discuss how the SOM comparison of the photometric and spectroscopic data could be used to improve GAMA representation with respect to KiDS-Bright. As a result, we may remove galaxies with ill-constrained photo-$z$ from the KiDS-Bright sample, as determined by direct comparisons with both GAMA itself and the extended spectroscopic sample (Espec) presented in Sec. 3.2.2.
As the left-most panel of Fig. 4 suggests, the simplest way to improve GAMA representation would be to lower the magnitude limit of KiDS-Bright (raise the flux threshold), as the majority of the incomplete SOM cells are at the faintest end. This would, however, affect the usefulness of the photometric sample for lensing studies by decreasing its effective depth, while it would not necessarily lead to direct improvement in photo-$z$ quality, as redshift estimates depend on the full 9-band information. We, therefore, compare such a more aggressive flux cut with SOM-based analysis of photo-$z$ quality, where we will consider both relative and absolute photo-$z$ errors as our test statistics. The former is defined as
$Dz\equiv\delta z/z_{\mathrm{spec}}$ (1)
(i.e. fractional difference, expressed in percent), where $\delta z\equiv z_{\mathrm{phot}}-z_{\mathrm{spec}}$. As photo-$z$ always have non-negligible uncertainties, including at very low redshift, this relative error $Dz$ is expected to diverge as $z\rightarrow 0$, but otherwise be stable for most of the redshift range. Our second statistic, the absolute photo-$z$ error, is defined as
@@
In Figs. 8-10 we show the distributions of the same galaxy properties as analyzed earlier in Sec. 4.1, but now comparing three cases in each panel: before any cuts to the KiDS-Bright dataset (filled tan), after the cuts (black), and additionally what is removed by the cuts (red). Fig. 8 applies to the SOM cells with the faintest sources being removed (item [1] above), Fig. 9 depicts the case where the cells with the largest relative photo-$z$ error $Dz$ are taken out (item [2]), while Fig. 10 illustrates the removal of the biggest absolute photo-$z$ errors ($\Delta z$, item [3]). Finally, Fig. 11 illustrates the joint condition [4] discussed above. For item [4], we also show the histograms for the objects discarded by the individual cuts that are combined, each of which culls about 4-5% of KiDS-Bright galaxies. Figures 9-11 are composed of two sets of panels each, those in the top based on using GAMA only as spec-$z$ calibration, while the bottom ones employ the entire Espec compilation for the cuts.
We can summarize the results as follows:
@@
    Removing the SOM cells at the faint end affects mostly galaxies that are redder, have larger stellar masses, and are at higher redshifts than the rest of the KiDS-Bright sample. This is expected for a flux-limited dataset, where the (observationally) faint end is dominated by intrinsically bright, red, massive galaxies, which on average are observed at higher redshifts than the sample mean.
@@
    Cleaning up the KiDS-Bright sample of the SOM cells that have the largest absolute photo-$z$ errors, $\Delta z$, also gives quite different outcomes than cases [1] and [2]. While still preferably galaxies at the fainter (observed) magnitudes are affected, and practically all the bluest objects are discarded this way, the distributions of the $g-r$ color, stellar mass, and photo-$z$ of the removed sources otherwise sample these parameters much more uniformly. In other words, requiring a fixed cut in absolute photo-$z$ error seems to remove galaxies equally as a function of color, stellar mass, and photo-$z$. It is however important to note that this is not equivalent to a sparse-sampling, as the distributions of the removed objects are very different from those of the entire input population.
@@
By comparing the upper and lower panels of Figs. 9-11 we can see some differences between the galaxies affected by the cuts when using GAMA versus Espec as calibration. This can indicate situations where GAMA does not provide sufficient calibration for KiDS-Bright color space and additional data, such as from Espec, should be used to purify the photometric sample. However, as discussed below, interchanging between the two spectroscopic datasets for cleanup has very little influence on the resulting photo-$z$ statistics.
Table 2: Statistics of photometric redshift performance for the KiDS-Bright sample cross-matched with GAMA-eq. The numbers in columns (2)-(7) refer to subsamples kept after applying the cuts in the first column.
@@
The effect of the KiDS-Bright sample cleanup on its photo-$z$ properties is quantified in Table 2. The photo-z statistics is calculated for the KiDS-Bright sample cross-matched with GAMA-eq. As expected, the average redshift of the sample decreases if a brighter flux cut is adopted, while it goes up if the galaxies with the largest relative photo-$z$ error $Dz$ (Eq. 1) are discarded, and it remains largely unchanged if the cleanup is based on the largest absolute errors $\Delta z$. The reduction in the mean photo-$z$ residuals (Eq. 2, $4^{\rm th}$ and $5^{\rm th}$ columns) is by a factor of a few or even more (in absolute terms) for redshift-based cuts while interestingly they stay almost intact for a more aggressive magnitude cut. Furthermore, each of the analyzed cuts lowers the scatter in photo-$z$ errors, both in terms of standard deviation and SMAD, when compared to the original sample case: the former is reduced by $\sim 6\%$ (when implementing cleanup in $Dz$ or $\Delta z$), whereas the latter decreases by $\sim 3\%$ (cut in $Dz$ based on the Espec calibration). It is worth noting that a larger relative reduction in SD than in SMAD means that the photo-$z$ residuals become more Gaussian after our proposed cuts, due to a reduction in outliers. Still, the fact that SD is still much larger than SMAD indicates that there are still considerable non-Gaussian wings and/or outlier populations in the photo-$z$ error distribution (see, e.g., the discussion in B21).
Finally, we note that the same general observations hold when we use more conservative selection cuts and remove twice as many galaxies (i.e. 20%) from the sample; see Appendix A. In fact, one would need to remove the faintest half of the dataset (cutting at $r\lesssim 19.4$ mag) to obtain only 10% reduction in photo-$z$ scatter, as we shortly discuss in Appendix B. Similarly, reducing the mean residuals to $\sim 0$ would require rejecting over 30% of the KiDS-Bright galaxies. Such aggressive cleanups would considerably affect the usefulness of the dataset, especially for GGL science. Indeed, as shown in e.g. B21, for this sample it is the actual number of galaxies used as lenses that drives the uncertainties of the GGL measurements, rather than the photo-$z$ errors. In other words, to improve the scientific usefulness of the KiDS-Bright dataset, it would be more desirable to increase its number density, rather than to reduce it.
@@ 5 Summary and conclusions
In this paper, we employed self-organizing maps (SOMs) to analyze how well the photometric KiDS-Bright galaxy sample, flux-limited at $r<20$ mag, is matched in color space with available overlapping spectroscopic data samples. This analysis is particularly important because applications of the KiDS-Bright sample often require robust photometric redshift estimates. Therefore, it is crucial to validate the performance of the current ANNz2-based photo-$z$ (B21) across various galaxy types and brightness ranges.
We first focused on the GAMA-Equatorial (GAMA-Eq) spec-$z$ sample, which was used by B21 as a training dataset in ANNz2 to derive KiDS-Bright photo-$z$. GAMA-Eq was originally designed to be highly complete spectroscopically up to $r<19.8$ in Petrosian magnitude, as derived from SDSS (Baldry et al., 2010). Its eventual completeness was, however, established to be 98% at $r<19.6$ using KiDS imaging (Driver et al., 2022), which is 0.4 mags brighter than the KiDS-Bright flux limit. Nonetheless, as GAMA includes galaxies with spec-$z$ measurements fainter than this limit (typically filler targets), B21 was able to train KiDS-Bright photo-$z$ beyond the GAMA completeness limit. Our SOM analysis shows that while GAMA-Eq completeness, relative to KiDS-Bright, quickly declines for magnitudes fainter than $r\sim 19.5$, there is no significant trend in relative incompleteness with multi-wavelength colors or stellar mass. This suggests that, despite missing many faint-end galaxies, GAMA-Eq does provide a good representation of the KiDS-Bright galaxy population over the color space.
Having compared the KiDS-Bright and GAMA in the color space using SOM, we next proposed a possible approach to further enhance the photo-$z$ quality of the former, by cleaning the target sample to force a better match with the overlapping spectroscopy. Here we extend beyond GAMA, by adding spec-$z$ from several other datasets overlapping with KiDS, as well as with external calibration fields dubbed ‘KiDZ’.
@@
We conclude that none of the cleanups that we analyzed led to significant improvement in the overall photo-$z$ quality of the KiDS-Bright sample. The SMAD values never reduce below 3% even after 10% reduction in the sample. On the one hand, this may indicate that the current KiDS-Bright photo-$z$ are already close to being optimal for the set of pass-bands that KiDS+VIKING provide.
@@
Our findings presented here can be further scrutinized in the near future by supplementing the spec-$z$ data we used here with additional redshifts. This could be done by either selecting galaxies from the most under-represented cells for targeted spectroscopic follow-up (similarly as done in the C3R2 projects, Masters et al. 2017) or by gradually supplementing these cells with spec-$z$ coming from large surveys. In this latter respect, DESI looks particularly promising, as it includes a Bright Galaxy Survey (Hahn et al., 2023), in which the faint sub-survey is obtaining redshifts for sources up to $r<20.175$. In our work, we include DESI Early Data Release galaxies (DESI Collaboration et al., 2023), but their overlap with our photometric sample is evidently too small to considerably influence our findings. We hope to be able to further test our approach with the forthcoming DESI Data Release 1 from the main survey. On a longer term, the 4-meter Multi-Object Spectroscopic Telescope (4MOST, de Jong et al., 2019) will undertake a number of surveys that should provide a good match to KiDS both in sky coverage and color space.
In this work, we focused on the KiDS DR4 Bright Sample, but our framework can be adapted to any galaxy dataset where such an analysis is relevant. In particular, we plan to employ the selection of foreground galaxies from the final data of KiDS DR5 (Wright et al., 2024), where a new version of the KiDS Bright Sample will be designed as the main dataset for GGL and clustering in KiDS 3$\times$2pt analyses.
Acknowledgements
@@
PJ and MB are supported by the Polish National Science Center through grant no. 2020/38/E/ST9/00395. MB and WAH are supported by the Polish National Science Center through grants no. 2018/30/E/ST9/00698, 2018/31/G/ST9/03388 and 2020/39/B/ST9/03494. AHW is supported by the Deutsches Zentrum für Luft- und Raumfahrt (DLR), made possible by the Bundesministerium für Wirtschaft und Klimaschutz, and acknowledges funding from the German Science Foundation DFG, via the Collaborative Research Center SFB1491 ”Cosmic Interacting Matters - From Source to Signal”. AD acknowledged support from ERC Consolidator Grant (No. 770935). CH acknowledges support from the European Research Council under grant number 647112, from the Max Planck Society and the Alexander von Humboldt Foundation in the framework of the Max Planck-Humboldt Research Award endowed by the Federal Ministry of Education and Research, and the UK Science and Technology Facilities Council (STFC) under grant ST/V000594/1. HHi is supported by a DFG Heisenberg grant (Hi 1495/5-1), the DFG Collaborative Research Center SFB1491, as well as an ERC Consolidator Grant (No. 770935). KK acknowledges support from the Royal Society and Imperial College. CM acknowledges support under the grant number PID2021-128338NB-I00 from the Spanish Ministry of Science, and from the European Research Council under grant agreement No. 770935. SJN is supported by the US National Science Foundation (NSF) through grant AST-2108402. JLvdB is supported by an ERC Consolidator Grant (No. 770935). MY acknowledges funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant agreement No. 101053992).
This work has made use of TOPCAT (Taylor, 2005) software, as well as of python (www.python.org), including the packages NumPy (van der Walt et al., 2011), SciPy (Jones et al., 2001–), AstroPy (Astropy Collaboration et al., 2022), Pandas (Wes McKinney, 2010), Seaborn (Waskom, 2021) and Matplotlib (Hunter, 2007).
@@
Author contributions: All authors contributed to the development and writing of this paper. The authorship list is given in two groups: the lead authors (PJ, MB, WAH, AHW), followed by an alphabetical group which includes those who have either made a significant contribution to the data products, or to the scientific analysis.

### v2

@@ 4.2 Selection of a clean sample using SOM
In the previous section, we quantified the relative completeness of the GAMA spectroscopic data with respect to the photometric KiDS-Bright as a function of several galaxy properties. We observed that the GAMA completeness quickly drops at the faintest and highest-$z$ end, while no significant deterioration is present with regard to colors or stellar mass. As GAMA was used to train the KiDS-Bright photo-$z$, here we discuss how the SOM comparison of the photometric and spectroscopic data could be used to improve GAMA representation with respect to KiDS-Bright. As a result, we may remove galaxies with ill-constrained photo-$z$ from the KiDS-Bright sample, as determined by direct comparisons with both GAMA itself and the extended spectroscopic sample (Espec) presented in Sect. 3.2.2.
As the left-most panel of Fig. 4 suggests, the simplest way to improve GAMA representation would be to lower the magnitude limit of KiDS-Bright (raise the flux threshold), as the majority of the incomplete SOM cells are at the faintest end. This would, however, affect the usefulness of the photometric sample for lensing studies by decreasing its effective depth, while it would not necessarily lead to direct improvement in photo-$z$ quality, as redshift estimates depend on the full nine-band information. We therefore compare such a more aggressive flux cut with SOM-based analysis of photo-$z$ quality, where we will consider both relative and absolute photo-$z$ errors as our test statistics. The former is defined as
$Dz\equiv\delta z/z_{\mathrm{spec}},$ (1)
namely, as a fractional difference, expressed in percent, where $\delta z\equiv z_{\mathrm{phot}}-z_{\mathrm{spec}}$. As photo-$z$ always have non-negligible uncertainties, including at very low redshift, this relative error $Dz$ is expected to diverge as $z\rightarrow 0$, but otherwise be stable for most of the redshift range. Our second statistic, the absolute photo-$z$ error, is defined as
@@
In Figs. 8-10 we show the distributions of the same galaxy properties as analyzed earlier in Sect. 4.1, but now comparing three cases in each panel: before any cuts to the KiDS-Bright dataset (filled tan), after the cuts (black), and additionally what is removed by the cuts (red). Fig. 8 applies to the SOM cells with the faintest sources being removed (item [1] above), Fig. 9 depicts the case where the cells with the largest relative photo-$z$ error $Dz$ are taken out (item [2]), while Fig. 10 illustrates the removal of the biggest absolute photo-$z$ errors ($\Delta z$, item [3]). Finally, Fig. 11 illustrates the joint condition [4] discussed above. For item [4], we also show the histograms for the objects discarded by the individual cuts that are combined, each of which culls about 4-5% of KiDS-Bright galaxies. Figures 9-11 are composed of two sets of panels each, those in the top based on using GAMA only as spec-$z$ calibration, while the bottom ones employ the entire Espec compilation for the cuts. We can summarize the results as follows:
@@
    Removing the SOM cells at the faint end affects mostly galaxies that are redder, have higher stellar masses, and are at higher redshifts than the rest of the KiDS-Bright sample. This is expected for a flux-limited dataset, where the (observationally) faint end is dominated by intrinsically bright, red, massive galaxies, which are typically observed at higher redshifts than the sample mean.
@@
    Cleaning up the KiDS-Bright sample of the SOM cells that have the largest absolute photo-$z$ errors, $\Delta z$, also gives quite different outcomes than cases [1] and [2]. While it is still preferably galaxies at the fainter (observed) magnitudes that are affected, and practically all the bluest objects are discarded this way, the distributions of the $g-r$ color, stellar mass, and photo-$z$ of the removed sources otherwise sample these parameters much more uniformly. In other words, requiring a fixed cut in absolute photo-$z$ error seems to remove galaxies equally as a function of color, stellar mass, and photo-$z$. It is however important to note that this is not equivalent to a sparse sampling, as the distributions of the removed objects are very different from those of the entire input population.
@@
By comparing the upper and lower panels of Figs. 9-11, we can see some differences between the galaxies affected by the cuts when using GAMA versus Espec as calibration. This can indicate situations where GAMA does not provide sufficient calibration for KiDS-Bright color space and additional data, such as from Espec, should be used to purify the photometric sample. However, as discussed below, interchanging between the two spectroscopic datasets for cleanup has very little influence on the resulting photo-$z$ statistics.
Table 2: Statistics of photometric redshift performance for the KiDS-Bright sample cross-matched with GAMA-eq. Numbers in columns (4)-(7) refer to subsamples kept after applying the cuts in the first column.
@@
The effect of the KiDS-Bright sample cleanup on its photo-$z$ properties is quantified in Table 2. As expected, the average redshift of the sample decreases if a brighter flux cut is adopted, while it goes up if the galaxies with the largest relative photo-$z$ error $Dz$ (Eq. 1) are discarded. It remains largely unchanged if the cleanup is based on the largest absolute errors $\Delta z$. The photo-z statistics are calculated for the KiDS-Bright sample cross-matched with GAMA-eq. The reduction in the mean photo-$z$ residuals (Eq. 2, columns 4-5) is by a factor of a few or even more (in absolute terms) for redshift-based cuts while interestingly they stay almost intact for a more aggressive magnitude cut. Furthermore, each of the analyzed cuts lowers the scatter in photo-$z$ errors, both in terms of standard deviation and SMAD, when compared to the original sample case: the former is reduced by $\sim 6\%$ (when implementing cleanup in $Dz$ or $\Delta z$), whereas the latter decreases by $\sim 3\%$ (cut in $Dz$ based on the Espec calibration). It is worth noting that a larger relative reduction in SD than in SMAD means that the photo-$z$ residuals become more Gaussian after our proposed cuts, due to a reduction in outliers. Still, the fact that SD is still much larger than SMAD indicates that there are still considerable non-Gaussian wings and/or outlier populations in the photo-$z$ error distribution (see, e.g., the discussion in B21).
Finally, we note that the same general observations hold when we use more conservative selection cuts and remove twice as many galaxies (i.e., 20%) from the sample; see Appendix A. In fact, one would need to remove the faintest half of the dataset (cutting at $r\lesssim 19.4$ mag) to obtain only 10% reduction in photo-$z$ scatter, as we shortly discuss in Appendix B. Similarly, reducing the mean residuals to $\sim 0$ would require rejecting over 30% of the KiDS-Bright galaxies. Such aggressive cleanups would considerably affect the usefulness of the dataset, especially for GGL science. Indeed, as shown in B21, for example, for this sample, it is the actual number of galaxies used as lenses that drives the uncertainties of the GGL measurements, rather than the photo-$z$ errors. In other words, to improve the scientific usefulness of the KiDS-Bright dataset, it would be more desirable to increase its number density, rather than to reduce it.
@@ 5 Summary and conclusions
In this paper, we employed self-organizing maps (SOMs) to analyze how well the photometric KiDS-Bright galaxy sample, flux-limited at $r<20$ mag, is matched in the color space with the available overlapping spectroscopic data samples. This analysis is particularly important because applications of the KiDS-Bright sample often require robust photometric redshift estimates. Therefore, it is crucial to validate the performance of the current ANNz2-based photo-$z$ (B21) across various galaxy types and brightness ranges.
We first focused on the GAMA-Equatorial (GAMA-Eq) spec-$z$ sample, which was used by B21 as a training dataset in ANNz2 to derive KiDS-Bright photo-$z$. GAMA-Eq was originally designed to be highly complete spectroscopically up to $r<19.8$ in Petrosian magnitude, as derived from SDSS (Baldry et al., 2010). Its eventual completeness was, however, established to be 98% at $r<19.6$ using KiDS imaging (Driver et al., 2022), which is 0.4 mags brighter than the KiDS-Bright flux limit. Nonetheless, as GAMA includes galaxies with spec-$z$ measurements fainter than this limit (typically filler targets), B21 was able to train KiDS-Bright photo-$z$ beyond the GAMA completeness limit. Our SOM analysis shows that while GAMA-Eq completeness, relative to KiDS-Bright, quickly declines for magnitudes fainter than $r\sim 19.5$, there is no significant trend in relative incompleteness with multiwavelength colors or stellar mass. This suggests that, despite missing many faint-end galaxies, GAMA-Eq does provide a good representation of the KiDS-Bright galaxy population over the color space.
Having compared the KiDS-Bright and GAMA in the color space using SOM, we next proposed a possible approach to further enhance the photo-$z$ quality of the former, by cleaning the target sample to force a better match with the overlapping spectroscopy. Here, we extend our study beyond GAMA, by adding spec-$z$ from several other datasets overlapping with KiDS, as well as with external calibration fields dubbed ”KiDZ.”
@@
We conclude that none of the cleanups that we analyzed led to significant improvement in the overall photo-$z$ quality of the KiDS-Bright sample. The SMAD values are never reduced to below 3%, even after 10% reduction in the sample. On the one hand, this may indicate that the current KiDS-Bright photo-$z$ are already close to being optimal for the set of passbands that KiDS+VIKING provide.
@@
Our findings presented here can be further scrutinized in the near future by supplementing the spec-$z$ data we used here with additional redshifts. This could be done by either selecting galaxies from the most under-represented cells for targeted spectroscopic follow-up (similarly as done in the C3R2 projects, Masters et al. 2017) or by gradually supplementing these cells with spec-$z$ coming from large surveys. In this latter respect, DESI looks particularly promising, as it includes a Bright Galaxy Survey (Hahn et al., 2023), in which the faint sub-survey is obtaining redshifts for sources up to $r<20.175$. In our work, we include DESI Early Data Release galaxies (Collaboration et al., 2024), but their overlap with our photometric sample is evidently too small to considerably influence our findings. We hope to be able to further test our approach with the forthcoming DESI Data Release 1 from the main survey. On a longer term, the 4-meter Multi-Object Spectroscopic Telescope (4MOST, de Jong et al., 2019) will undertake a number of surveys that should provide a good match to KiDS both in sky coverage and color space.
In this work, we focused on the KiDS DR4 Bright Sample, but our framework can be adapted to any galaxy dataset where such an analysis is relevant. In particular, we plan to employ the selection of foreground galaxies from the final data of KiDS DR5 (Wright et al., 2024), where a new version of the KiDS Bright Sample will be designed as the main dataset for GGL and clustering in KiDS 3$\times$2pt analyses. Furthermore, our methodology will be relevant for forthcoming deep wide-area surveys such as from Euclid or LSST. While a fraction of galaxies detected by these surveys will have been covered by spectroscopy from DESI, 4MOST, or Euclid itself, on their full area the most complete selection of foreground galaxies for clustering and galaxy-galaxy lensing may still be possible only from photometry. It will be then natural to apply similar approaches to photo-z derivation for, for instance, KiDS-Bright (i.e., by training some ML models), which will equally need quantification of the match between spectroscopic training sets and target photometrically selected samples. Our framework presented here could be practically directly employed once these new data become available. We can then follow up with an appropriate clean-up of the photometric datasets and observational proposals to obtain spectroscopy for the least complete SOM cells.
Acknowledgements.
@@
We thank our anonymous referee for their positive feedback and useful comments. PJ and MB are supported by the Polish National Science Center through grant no. 2020/38/E/ST9/00395. MB and WAH are supported by the Polish National Science Center through grants no. 2018/30/E/ST9/00698, 2018/31/G/ST9/03388 and 2020/39/B/ST9/03494. AHW is supported by the Deutsches Zentrum für Luft- und Raumfahrt (DLR), made possible by the Bundesministerium für Wirtschaft und Klimaschutz, and acknowledges funding from the German Science Foundation DFG, via the Collaborative Research Center SFB1491 ”Cosmic Interacting Matters - From Source to Signal”. AD acknowledged support from ERC Consolidator Grant (No. 770935). This publication is part of the project “A rising tide: Galaxy intrinsic alignments as a new probe of cosmology and galaxy evolution” (with project number VI.Vidi.203.011) of the Talent programme Vidi which is (partly) financed by the Dutch Research Council (NWO). This work is also part of the Delta ITP consortium, a program of the Netherlands Organisation for Scientific Research (NWO) that is funded by the Dutch Ministry of Education, Culture and Science (OCW). CH acknowledges support from the European Research Council under grant number 647112, from the Max Planck Society and the Alexander von Humboldt Foundation in the framework of the Max Planck-Humboldt Research Award endowed by the Federal Ministry of Education and Research, and the UK Science and Technology Facilities Council (STFC) under grant ST/V000594/1. HHi is supported by a DFG Heisenberg grant (Hi 1495/5-1), the DFG Collaborative Research Center SFB1491, as well as an ERC Consolidator Grant (No. 770935). KK acknowledges support from the Royal Society and Imperial College. CM acknowledges support under the grant number PID2021-128338NB-I00 from the Spanish Ministry of Science, and from the European Research Council under grant agreement No. 770935. SJN is supported by the US National Science Foundation (NSF) through grant AST-2108402. JLvdB is supported by an ERC Consolidator Grant (No. 770935). ZY acknowledges support from the Max Planck Society and the Alexander von Humboldt Foundation in the framework of the Max Planck-Humboldt Research Award endowed by the Federal Ministry of Education and Research (Germany). MY acknowledges funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant Agreement No. 101053992). For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising from this submission. This work has made use of TOPCAT (Taylor, 2005) software, as well as of python (www.python.org), including the packages NumPy (van der Walt et al., 2011), SciPy (Jones et al., 2001–), AstroPy (Astropy Collaboration et al., 2022), Pandas (Wes McKinney, 2010), Seaborn (Waskom, 2021) and Matplotlib (Hunter, 2007).
@@
Author contributions: All authors contributed to the development and writing of this paper. The authorship list is given in two groups: the lead authors (PJ, MB, WAH, AHW), followed by an alphabetical group which includes those who have either made a significant contribution to the data products or to the scientific analysis.